Please help me solve this problem. 5y-8=3y+12, 5y-8+8=3y+12-8
8y=4
y=0.5

Answer by stanbon(75887)

Please help me solve this problem.
5y-8=3y+12
5y-8+8 = 3y+12+8
5y = 3y+20
5y-3y = 3y-3y+20
2y = 20
y = 10